# On the numerically and cohomologically trivial automorphisms of elliptic surfaces II: $\chi(S)>0$

### v1

@@ Abstract.
In this second part we study first the group $\mathrm{Aut}_{\mathbb{Q}}(S)$ of numerically trivial automorphisms of a properly elliptic surface $S$, that is, of a minimal surface with Kodaira dimension $\kappa(S)=1$, in the case $\chi(S)\geq 1$. Our first surprising result is that, against what has been believed for over 40 years, we have nontrivial such groups for $p_{g}(S)>0$. Indeed, we show even that there is no absolute upper bound for their cardinalities $|\mathrm{Aut}_{\mathbb{Q}}(S)|$. At any rate, we give explicit and almost sharp upper bounds for $|\mathrm{Aut}_{\mathbb{Q}}(S)|$ in terms of the numerical invariants of $S$, as $\chi(S)$, the irregularity $q(S)$, and the bigenus $P_{2}(S)$. Moreover, we come quite close to a complete description of the possible groups $\mathrm{Aut}_{\mathbb{Q}}(S)$, and we give an effective criterion for surfaces to have trivial $\mathrm{Aut}_{\mathbb{Q}}(S)$.
@@ 1. Introduction
Hence the major problems to be considered are: to determine when these two special quotient groups may be nontrivial, to give explicit upper bounds for their cardinality in terms of the invariants of $X$, and to try to establish sharpness of these upper bounds by exhibiting appropriate examples.
@@
This paper is dedicated to the remaining case of minimal surfaces of Kodaira dimension $1$ with $\chi(S)>0$ (the case $\chi(S)=0$, characterized by the fact that the pluricanonical elliptic fibration has all the fibres multiples of a smooth elliptic curve, was treated in Part I): here $\mathrm{Aut}^{0}(S)$ is trivial.
The main purpose of this article is to revive the interest for this special but very subtle topic of research, which, like the Arab Phoenix, has a rebirth from its own ashes. Indeed, for over 40 years it was believed that these groups are trivial for $p_{g}(S)>0$ (Theorem 4.5 of [Pe80], confirmed by Cor. 0.3 of [Cai09]).
Recent results of Dolgachev–Martin [DM22] for the case where $S$ is rational of any characteristic (which we shall use) go also in this direction.
Our results, which will be more amply illustrated later on, are rather complete for the case of the group $\mathrm{Aut}_{\mathbb{Q}}(S)$ of numerically trivial automorphisms.
For the intriguing case of the group $\mathrm{Aut}_{\mathbb{Z}}(S)$ of cohomologically trivial automorphisms we establish again, with very hard work, the existence (as in part I) of nontrivial groups of order $2,3$: but we cannot establish an absolute upper bound for $|\mathrm{Aut}_{\mathbb{Z}}(S)|$ except under special assumptions, $p_{g}(S)=0$, or if $S$ is isotrivially fibred.
@@
To explain the difficulty of the question, the problem is that, even if one shows trivial action on $H_{1}(S,\mathbb{Z})=H^{3}(S,\mathbb{Z})$, there remains to handle the case of $H^{2}(S,\mathbb{Z})$, which has a non canonical splitting as a free part plus a torsion subgroup, canonically isomorphic to the torsion of $H_{1}(S,\mathbb{Z})$. Hence the difficulty is to find a system of explicit generators for $H^{2}(S,\mathbb{Z})$, and this is quite hard in general, and almost impossible in the case where there is nontrivial torsion and $p_{g}(S)>0$.
Let us pass now to describe the main ideas and tools used in this paper.
@@
The most algebraic point of view is to look at the function field extension $\mathbb{C}(S)\supset\mathbb{C}(B)$ as providing a curve ${\mathcal{C}}$ of genus $1$ (the fibre of $f$ over the generic point of $B$) defined over the non algebraically closed field ${\mathcal{K}}:=\mathbb{C}(B)$. This point of view is essential to have a glimpse at $\mathrm{Aut}(S)$ in a simple algebraic way (observe indeed that in our case the group $\mathrm{Bir}(S)$ of birational self maps of $S$ equals $\mathrm{Aut}(S)$).
On the other hand the more topological point of view is to study the corresponding relative minimal model, that is, the fibration $f\colon S\rightarrow B$ and its singular fibres, and this will be crucial for checking when automorphisms are numerically trivial (respectively cohomologically trivial) or not.
The algebraic approach tells us immediately that, if we look at the subgroup $\mathrm{Aut}_{{\mathcal{K}}}({\mathcal{C}})$ of the automorphisms which act as the identity on the base field ${\mathcal{K}}$, these embed in $\mathrm{Aut}_{\bar{{\mathcal{K}}}}({\mathcal{C}}\otimes\bar{{\mathcal{K}}})$, where ${\mathcal{E}}:={\mathcal{C}}\otimes\bar{{\mathcal{K}}}$ is an elliptic curve over an algebraically closed field of characteristic zero,*** Indeed, it suffices to take a finite field extension ${\mathcal{K}}^{\prime}\supset{\mathcal{K}}$ to get an elliptic curve hence
@@
($r=4$ for the Gaussian = harmonic elliptic curve, $r=6$ for the Fermat = equianharmonic elliptic curve). Here $\mu_{r}$ is the group of $r$th roots of $1$ in $\mathbb{C}$, and ${\mathcal{E}}$ acts by translations via the isomorphism ${\mathcal{E}}\cong\operatorname{Pic}^{1}({\mathcal{E}})\cong\operatorname{Pic}^{0}({\mathcal{E}})$ provided by the choice of a point $0\in{\mathcal{E}}$ ($\operatorname{Pic}^{d}({\mathcal{E}})$ is as usual the set of linear equivalence clases of divisors of degree $d$).
If we go back to ${\mathcal{C}}$, there is no longer (as we were taught by Kodaira [Kod63] and Shafarevich [Shaf-65, Chapter VII]) an isomorphism ${\mathcal{C}}\cong\operatorname{Pic}^{0}({\mathcal{C}})$ as curves over the field ${\mathcal{K}}$ if there is no ${\mathcal{K}}$-rational point on ${\mathcal{C}}$, but the isomorphism ${\mathcal{C}}\cong\operatorname{Pic}^{1}({\mathcal{C}})$ holds true (by Riemann Roch and by descent).
Then we denote as usual $\mathrm{Jac}({\mathcal{C}}):=\operatorname{Pic}^{0}({\mathcal{C}})$, and ${\mathcal{C}}$ is a principal homogeneous space over $\mathrm{Jac}({\mathcal{C}})$ ([Shaf-65], Chapter 2 of [Dolg81]), in particular, the Mordell–Weil group, the group of ${\mathcal{K}}$-rational points of $\mathrm{Jac}({\mathcal{C}})$ acts on ${\mathcal{C}}$.
The torsion subgroup of the Mordell–Weil group will be a major source for the construction of numerically trivial automorphisms.
The geometrical counterpart of $\mathrm{Jac}({\mathcal{C}})$ is the Jacobian surface $J(S)$, a suitable compactification of ${\mathcal{R}}^{1}f_{*}\mathcal{O}_{S}/{\mathcal{R}}^{1}f_{*}\mathbb{Z}$, with the property that the singular fibres of $J(S)$ are the same as the singular fibres $F$ of $S$ when $F$ is not multiple, while they are isogenous to $F_{\mathrm{red}}$ when $F$ is multiple. In this way, even if the global topological behaviour of $J(S)$ differs from that of $S$, at least the topological nature of the fibres is the same.
@@
here the left hand side group by (1.1) contains the Mordell–Weil group as a subgroup of index $r\leq 6$
@@
The strategy is of course to give upper bounds for $|\mathrm{Aut}(S)|$ via upper bounds for the cardinalities $|\mathrm{Aut}(S)|_{B}|$ of the image group (here an essential role is given by the analysis of the singular fibres) and $|\mathrm{Aut}_{B}(S)|$ of the group of automorphisms acting trivially on the base curve $B$.
Here, the analysis bifurcates: if the fibration is isotrivial, that is, birational to a quotient $(C\times E)/G$ (where the action of $G$ is no longer assumed to be free), then we use many of the tools developed in Part I, but now enhanced by the presence of reducible fibres.
Whereas, if the fibration is not isotrivial (that is, the smooth fibres are not all isomorphic), then the Mordell–Weil group plays a dominant role: since then, in view of the finiteness of $\mathrm{Aut}_{\mathbb{Q}}(S)$, we have to study when elements of the torsion subgroup $\mathrm{MW}(J(S))_{\mathrm{tor}}$ are numerically trivial.
@@
And when $J(S)$ is not rational, we can view it as a pull back from an elliptic modular surface, and this paves the way to exhibiting large groups of numerically trivial automorphisms.
@@
1.1. An outline of the main results of the paper
Let $f\colon S\to B$ denote a properly elliptic surface with $\chi(S)>0$. Our main theorems show that the group $\mathrm{Aut}_{\mathbb{Q}}(S)$ of numerically trivial automorphism can be arbitrarily large, but in a controlled way. We do this by controlling the induced action of $\mathrm{Aut}_{\mathbb{Q}}(S)$ on the base $B$ of the fibration and then by relating $\mathrm{Aut}_{\mathbb{Q}}(S)$ to the torsion part $\mathrm{MW}(J(S))_{\mathrm{tor}}$ of the Mordell–Weil group of the relative Jacobian $J(S)$.
The behavior of $\mathrm{Aut}_{\mathbb{Q}}(S)$ differs a lot according to whether $p_{g}(S)$ is positive or not. In the case where $p_{g}(S)>0$, or, equivalently, where $J(S)$ is not rational, $\mathrm{Aut}_{\mathbb{Q}}(S)$ preserves each fibre and induces translations on smooth fibres, thus comes from a subgroup of $\mathrm{MW}(J(S))_{\mathrm{tor}}$, as is illustrated in Lemma 4.4. The upshot is then that, via our main construction of examples (Construction 7.5), we give a flexible sufficient condition for a subgroup $G\subset\mathrm{MW}(J(S))_{\mathrm{tor}}$ to be realized as a subgroup of $\mathrm{Aut}_{\mathbb{Q}}(S)$.
Theorem 1.1.
@@ 1.1. An outline of the main results of the paper
(ii) Conversely, for any finite 2-generated abelian group $G$, there is a properly elliptic surface $S$ such that $\mathrm{Aut}_{\mathbb{Q}}(S)\supseteq G$. Equality $\mathrm{Aut}_{\mathbb{Q}}(S)=G$ can be ensured if we have an isomorphism $G\cong\mathbb{Z}/N\mathbb{Z}\times\mathbb{Z}/N^{\prime}\mathbb{Z}$ such that at every prime $p$, the $p$-adic valuations satisfy $v_{p}(N)=v_{p}(N^{\prime})$ or $\max(v_{p}(N),v_{p}(N^{\prime}))\leq 1$.
(iii) $\mathrm{Aut}_{\mathbb{Q}}(S)$ is trivial if there is an additive fibre or if all multiple fibres of $S\to B$ have smooth support; in particular, this holds true if there are no multiple fibres, e.g. if the fibration admits a section, that is, $S$ is a jacobian elliptic surface.
@@
By bounding the Mordell–Weil group of the elliptic modular surfaces $X(N)$ and $X_{1}(N)$ associated to the congruence modular groups $\Gamma(N)$ and $\Gamma_{1}(N)$, we obtain an upper bound for $|\mathrm{Aut}_{\mathbb{Q}}(S)|$ in the case $p_{g}(S)>0$. There are even some arithmetic relation between the prime exponents of $\mathrm{Aut}_{\mathbb{Q}}(S)$ and the Euler characteristic $\chi(S)$.
Theorem 1.2.
@@
Details about the bounds in Theorem 1.2 can be found in Sections 8.5, 8.6. We emphasize that the bound in (1.3) also holds if $p_{g}(S)=0$ outside a few cases, see Remark 6.14. This can also be deduced from (and improved by) our main theorem in the case $p_{g}=0$, see also Theorem 3.16 for part (ii).
Theorem 1.3.
@@
    conversely, for any $s\in\mathbb{N}$, there is $S$ with $P_{2}(S)=s-1$, $|\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}|=s$ and $|\mathrm{Aut}_{\mathbb{Q}}(S)|=4s$.
@@
We should point out that, for ease of presentation, the theorem is largely simplified in the sense that for many isotrivial fibrations the bounds from (i) hold, cf. Proposition 6.1.
Cohomologically trivial automorphisms are much rarer than numerically trivial ones.
In contrast to Theorem 1.3, we have a uniform bound for $\mathrm{Aut}_{\mathbb{Z}}(S)$ in the case $p_{g}=0.$ The actual existence of properly elliptic surfaces with nontrivial $\mathrm{Aut}_{\mathbb{Z}}(S)$ is also surprising in view of the previous statements in the literature (cf. [Pe80, Theorem 4.5]).
Theorem 1.4.

### v2

@@ Abstract.
In this second part we study first the group $\mathrm{Aut}_{\mathbb{Q}}(S)$ of numerically trivial automorphisms of an algebraic properly elliptic surface $S$, that is, of a minimal algebraic surface with Kodaira dimension $\kappa(S)=1$, in the case $\chi(S)\geq 1$. Our first surprising result is that, against what has been believed for over 40 years, there exist nontrivial such groups for $p_{g}(S)>0$. Indeed, we show even that there is no absolute upper bound for their cardinalities $|\mathrm{Aut}_{\mathbb{Q}}(S)|$. At any rate, we give explicit and nearly optimal upper bounds for $|\mathrm{Aut}_{\mathbb{Q}}(S)|$ in terms of the numerical invariants of $S$, as $\chi(S)$, or the irregularity $q(S)$, or the bigenus $P_{2}(S)$. Moreover, we come quite close to a complete description of the possible groups $\mathrm{Aut}_{\mathbb{Q}}(S)$ as 2-generated finite abelian groups and we give effective criteria for surfaces to have trivial $\mathrm{Aut}_{\mathbb{Q}}(S)$.
@@ 1. Introduction
Hence the major problems to be considered are: to determine when this quotient group or $\mathrm{Aut}_{\mathbb{Z}}(X)/\mathrm{Aut}^{0}(X)$ may be nontrivial, to give explicit upper bounds for their cardinality in terms of the invariants of $X$, and to try to establish sharpness of these upper bounds by exhibiting appropriate examples.
@@
This paper is dedicated to the remaining case of minimal surfaces of Kodaira dimension $1$ with $\chi(S)>0$: they are automatically Kähler, and for them $\mathrm{Aut}^{0}(S)$ is trivial; we shall deal here with the case of algebraic surfaces, leaving the case of non algebraic surfaces for a future paper.
The main purpose of this article is to revive the interest for this very subtle topic of research, which, like the Arab Phoenix, has a rebirth from its own ashes. Indeed, for over 40 years it was believed that these groups are trivial for $p_{g}(S)>0$ (Theorem 4.5 of [Pe80], confirmed by Cor. 0.3 of [Cai09]).
Recent results of Dolgachev–Martin [DM22] for the case where $S$ is a rational surface of any characteristic (which we shall use) go also in the direction of investigating numerically trivial automorphisms (these are automatically cohomologically trivial if the surfaces are rational).
Our results, which will be more amply illustrated later on in 1.3, are rather complete for the case of the group $\mathrm{Aut}_{\mathbb{Q}}(S)$ of numerically trivial automorphisms.
For the intriguing case of the group $\mathrm{Aut}_{\mathbb{Z}}(S)$ of cohomologically trivial automorphisms we establish again, with very hard work, the existence (as in part I) of nontrivial groups of order $2,3$: but we cannot establish an absolute upper bound for $|\mathrm{Aut}_{\mathbb{Z}}(S)|$ except under special assumptions, for instance if $p_{g}(S)=0$, or if $S$ is isotrivially fibred.
@@
To explain the difficulty of the question, the problem is that, even if one shows that an automorphism has trivial action on $H_{1}(S,\mathbb{Z})=H^{3}(S,\mathbb{Z})$, there remains to handle the case of $H^{2}(S,\mathbb{Z})$, which has a non canonical splitting as a free part plus a torsion subgroup, canonically isomorphic to the torsion of $H_{1}(S,\mathbb{Z})$. Hence the difficulty is to find a system of explicit generators for $H^{2}(S,\mathbb{Z})$ in terms of which we can then describe the action of automorphisms; this is quite hard in general, and almost impossible in the case where there is nontrivial torsion and $p_{g}(S)>0$.
1.1. Main ideas and tools: two allied points of view
@@
The most algebraic point of view is to look at the function field extension $\mathbb{C}(S)\supset\mathbb{C}(B)$ as providing a curve ${\mathcal{C}}$ of genus $1$ (the fibre of $f$ over the generic point of $B$) defined over the non algebraically closed field ${\mathcal{K}}:=\mathbb{C}(B)$. This point of view is essential to have a simple algebraic picture of $\mathrm{Aut}(S)$: since in our case the group $\mathrm{Aut}(S)$ equals the group $\mathrm{Bir}(S)$ of birational self maps of $S$, and $\mathrm{Bir}(S)=Aut(\mathbb{C}(S))=Aut({\mathcal{C}})$.
On the other hand the study of the topological properties of automorphisms is based on the relative minimal model $f\colon S\rightarrow B$, that is, on the action on the fibration and on its singular fibres; here there is a crucial difference for checking when automorphisms are numerically trivial, respectively cohomologically trivial.
The algebraic approach tells us immediately that, if we look at the subgroup $\mathrm{Aut}_{{\mathcal{K}}}({\mathcal{C}})$ of the automorphisms which act as the identity on the base field ${\mathcal{K}}$, these embed in $\mathrm{Aut}_{\bar{{\mathcal{K}}}}({\mathcal{C}}\otimes\bar{{\mathcal{K}}})$, where ${\mathcal{E}}:={\mathcal{C}}\otimes\bar{{\mathcal{K}}}$ is an elliptic curve over an algebraically closed field of characteristic zero,*** Indeed, it suffices to take a finite field extension ${\mathcal{K}}^{\prime}\supset{\mathcal{K}}$ to get an elliptic curve. hence
@@
($r=4$ for the Gaussian = harmonic elliptic curve, $r=6$ for the Fermat = equianharmonic elliptic curve). Here $\mu_{r}$ is the group of $r$th roots of $1$ in $\mathbb{C}$, and ${\mathcal{E}}$ acts by translations via the isomorphism ${\mathcal{E}}\cong\operatorname{Pic}^{1}({\mathcal{E}})\cong\operatorname{Pic}^{0}({\mathcal{E}})$ provided by the choice of a point $0\in{\mathcal{E}}$ ($\operatorname{Pic}^{d}({\mathcal{E}})$ is as usual the set of linear equivalence classes of divisors of degree $d$).
If we go back to ${\mathcal{C}}$, there is no longer (as we were taught by Kodaira [Kod63] and Shafarevich [Shaf65, Chapter VII]) an isomorphism ${\mathcal{C}}\cong\operatorname{Pic}^{0}({\mathcal{C}})$ as curves over the field ${\mathcal{K}}$ if there is no ${\mathcal{K}}$-rational point on ${\mathcal{C}}$, but the isomorphism ${\mathcal{C}}\cong\operatorname{Pic}^{1}({\mathcal{C}})$ holds true (by Riemann Roch and by descent).
Then we denote as usual $\mathrm{Jac}({\mathcal{C}}):=\operatorname{Pic}^{0}({\mathcal{C}})$, and ${\mathcal{C}}$ is a principal homogeneous space over $\mathrm{Jac}({\mathcal{C}})$ ([Shaf65], Chapter 2 of [Dolg81]), in particular, the Mordell–Weil group, the group of ${\mathcal{K}}$-rational points of $\mathrm{Jac}({\mathcal{C}})$ acts on ${\mathcal{C}}$.
The torsion subgroup of the Mordell–Weil group will be a major source for the construction of numerically trivial automorphisms (see especially Section 7).
The geometrical counterpart of $\mathrm{Jac}({\mathcal{C}})$ is the Jacobian surface $J(S)$, a suitable compactification of the sheaf of groups ${\mathcal{R}}^{1}f_{*}\mathcal{O}_{S}/{\mathcal{R}}^{1}f_{*}\mathbb{Z}$ (whose stalks have a connected component of the identity which is either an elliptic curve, or $\mathbb{C}^{*}$, or $\mathbb{C}$: in the last case one says that the fibre is of additive type).
$J(S)$ has the property that the singular fibres of $J(S)$ are the same as the singular fibres $F$ of $S$ when $F$ is not multiple, while they are isogenous to $F_{\mathrm{red}}$ when $F$ is multiple. In this way, even if the global topological behaviour of $J(S)$ differs from that of $S$, at least the topological nature of the fibres is the same.
@@
where the left hand side group by (1.1) contains the Mordell–Weil group as a subgroup of index $r\leq 6$.
@@
1.2. Main strategy: a dichotomy of cases
Based on (1.2), the strategy is of course to give upper bounds for $|\mathrm{Aut}(S)|$ via upper bounds for the cardinalities $|\mathrm{Aut}(S)|_{B}|$ of the image group (here an essential role is given by the analysis of the singular fibres) and $|\mathrm{Aut}_{B}(S)|$ of the group of automorphisms acting trivially on the base curve $B$. Obviously this approach can also be applied to compute $|\mathrm{Aut}_{\mathbb{Q}}(S)|$ and $|\mathrm{Aut}_{\mathbb{Z}}(S)|$.
Here, the analysis bifurcates: if the fibration is isotrivial, that is, birational to a quotient $(C\times E)/G$ (where the action of $G$ is no longer assumed to be free), then we use many of the tools developed in Part I, but now enhanced (as we shall see) by the presence of reducible fibres, which have to be of additive type.
Whereas, if the fibration is not isotrivial (that is, the smooth fibres are not all isomorphic), then the Mordell–Weil group plays a dominant role: since then, in view of the finiteness of $\mathrm{Aut}_{\mathbb{Q}}(S)$, we have to study when elements of the torsion subgroup $\mathrm{MW}(J(S))_{\mathrm{tor}}$ yield numerically trivial automorphisms.
@@ 1.2. Main strategy: a dichotomy of cases
And when $J(S)$ is not rational, the method of viewing it as a pull back from an elliptic modular surface paves the way to exhibiting large groups of numerically trivial automorphisms (in Section 7).
@@
1.3. Main results for numerically trivial automorphisms
Let $f\colon S\to B$ denote a properly elliptic algebraic surface with $\chi(S)>0$. Our main theorems show that the group $\mathrm{Aut}_{\mathbb{Q}}(S)$ of numerically trivial automorphisms can be arbitrarily large, but in a controlled way. We do this by controlling the induced action of $\mathrm{Aut}_{\mathbb{Q}}(S)$ on the base $B$ of the fibration and then by relating $\mathrm{Aut}_{\mathbb{Q}}(S)$ to the torsion part $\mathrm{MW}(J(S))_{\mathrm{tor}}$ of the Mordell–Weil group of the relative Jacobian $J(S)$.
The behaviour of $\mathrm{Aut}_{\mathbb{Q}}(S)$ differs a lot according to whether $p_{g}(S)$ is positive or not. In the case where $p_{g}(S)>0$, or, equivalently, where $J(S)$ is not rational, $\mathrm{Aut}_{\mathbb{Q}}(S)$ preserves each fibre and induces translations on smooth fibres, thus comes from a subgroup of $\mathrm{MW}(J(S))_{\mathrm{tor}}$, as is illustrated in Lemma 4.4. The upshot is then that, via our main construction of examples (Construction 7.5), we give a flexible sufficient condition for a subgroup $G\subset\mathrm{MW}(J(S))_{\mathrm{tor}}$ to be realized as a subgroup of $\mathrm{Aut}_{\mathbb{Q}}(S)$.
Theorem 1.1 (Groups for positive geometric genus).
@@ 1.3. Main results for numerically trivial automorphisms
(ii) Conversely, for any finite 2-generated abelian group $G$, there is a properly elliptic surface $S$ such that $\mathrm{Aut}_{\mathbb{Q}}(S)\supseteq G$. Equality $\mathrm{Aut}_{\mathbb{Q}}(S)=G$ can be ensured if we have an isomorphism $G\cong\mathbb{Z}/N\mathbb{Z}\times\mathbb{Z}/N^{\prime}\mathbb{Z}$ such that, at every prime $p$, the $p$-adic valuations satisfy $|v_{p}(N)-v_{p}(N^{\prime})|\leq 1$.
(iii) $\mathrm{Aut}_{\mathbb{Q}}(S)$ is trivial if there is a fibre of additive type or if all multiple fibres of $S\to B$ have smooth support; in particular, this holds true if there are no multiple fibres, e.g. if the fibration admits a section, that is, $S\cong J(S)$.
@@
By bounding the Mordell–Weil group of the elliptic modular surfaces $X(N)$ and $X_{1}(N)$ associated to the congruence modular groups $\Gamma(N)$ and $\Gamma_{1}(N)$, we obtain an upper bound for $|\mathrm{Aut}_{\mathbb{Q}}(S)|$ in the case $p_{g}(S)>0$. There are even some arithmetic relations between the prime exponents of $\mathrm{Aut}_{\mathbb{Q}}(S)$ and the Euler characteristic $\chi(S)$.
Theorem 1.2 (Bounds for $\mathrm{Aut}_{\mathbb{Q}}$ in positive geometric genus).
@@
Details about the bounds in Theorem 1.2 can be found in Subsections 8.5, 8.6. We emphasize that the bound in (1.3) also holds if $p_{g}(S)=0$, with exception of a few cases, see Remark 6.14. This can also be deduced from (and improved by) our main theorem in the case $p_{g}=0$, see also Theorem 3.16 for part (ii).
Theorem 1.3 (Bounds for $\mathrm{Aut}_{\mathbb{Q}}$ for geometric genus zero).
@@
    conversely, for any $s\in\mathbb{N}$, there is an elliptic surface $S$ with $P_{2}(S)=s-1$, $|\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}|=s$ and $|\mathrm{Aut}_{\mathbb{Q}}(S)|=4s$.
@@
We should point out that, for ease of presentation, the above theorem is largely simplified, in the sense that for many isotrivial fibrations the bounds from (i) hold, cf. Proposition 6.1.
1.4. Main results for cohomologically trivial automorphisms
Cohomologically trivial automorphisms are much rarer than numerically trivial ones. In contrast to Theorem 1.3, we have a uniform bound for $\mathrm{Aut}_{\mathbb{Z}}(S)$ in the case $p_{g}=0.$ The actual existence of properly elliptic surfaces with nontrivial $\mathrm{Aut}_{\mathbb{Z}}(S)$ is also surprising in view of the previous statements in the literature (cf. [Pe80, Theorem 4.5]).
Theorem 1.4 (Bounds for $\mathrm{Aut}_{\mathbb{Z}}$ for genus zero).